OpenStudy (superhelp101):

Using complete sentences, explain how to find the equation of the line, in standard form and slope–intercept form, passing through (–1, 5) and (1, 9). Also compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope–intercept form.

OpenStudy (anonymous):

ok we need a point and the slope Let's find the slope: y2-y2/x2-x1 9-5/1+1 m = 2 point slope formula: y - y1 = m(x-x1) y - 9 = 2(x - 1) y = 9 + 2x - 2 y = 2x + 7

OpenStudy (anonymous):

Let me know if you need clarification

OpenStudy (superhelp101):

I understand. What's the standard form?

OpenStudy (anonymous):

And the benefits: slope intercept form shows the y-intercept point-slope form shows a point on the graph

OpenStudy (anonymous):

y - 9 = 2(x-1) is in point-slope form and y = 2x + 7 is slope-intercept form

OpenStudy (superhelp101):

but it asks for standard and slope-intercept

OpenStudy (superhelp101):

is that like ax+by=c ?

OpenStudy (anonymous):

O didn't catch that. It's ax + by + c = 0 So that would be 2x - y + 7 = 0

OpenStudy (anonymous):

benefits in writing in standard form: you can clearly see the values of a b and c which can help in certain cases

OpenStudy (superhelp101):

oh okay! thank you :))

OpenStudy (anonymous):

no problem at all

OpenStudy (superhelp101):

you're the best